In calcite(CaCO3) the Ca2+ ion has a coordination number 6. Using the appropriate Pauling rule determine the ion environment around each O2- ion.
Answer #1

In calcite Ca+2 forms Hexagonal lattice occupy corners,face centres and edges.Each Ca2+ is sorrounded by 6 neighbors.

According to Pauling rule

Electrostatic Valency( EV)= Charge on ion/ Coordination number

In case of Ca+2

charge= 2 and CN= 6

Hence

EV= 2/6= 1/3

In case of Carbonate ion

1.Carbon charge = +4 and Oxygen charge is -2

2. Each C+4 ion is sorrounded with three O-2.

Hence , EV of C = 4/3

3. Three Oxygens each contribute 4/3 charge to the C+4 ion, So the charge on the carbon is balanced. But, each Oxide ion has 2/3 of residual charge that it has not used.

Thus, a carbonate (CO3-2) structural group with triangular planar structure is formed

To avoid confusion each ion environment explained in separate points.